Convex and concave mirror have the same radii of curvature R.The distance between the mirrors is 2R.At what point on the common optical axis of the mirror should a point source of light A be placed for the ray to covarage at point A after being reflected first on convex and then on concave mirror?

Grade:11

1 Answers

harini desiraju
13 years ago

Ans:R

Sol'n:let the source be at a distance X from concave mirror.Then for convex mirror,

(1/2R-X)-(1/V)=2/R

from this we get, V=R(2R-X)/2X-3R

for concave mirror,we have 

[1/(2R+V)]+(1/X)=2/R

substituting the value of V and furthur solving for X we get X=R.